OpenStudy (anonymous):

Clara invests $5,000 in an account that pays 6.25% interest per year. After how many years will her investment be worth $10,000? -12.6 -11.4 -13.7 -14.1

hero (hero):

Hint: Use compound interest formula

OpenStudy (anonymous):

i tired but my awnser is irrevelent to my choices... so idk i need sum1 2 break it down for me, im new to this subject

OpenStudy (anonymous):

y dont u try to explain it to me bkuz clearly his awnser didnt fit

OpenStudy (anonymous):

-12.6 -11.4 -13.7 -14.1

OpenStudy (anonymous):

The compound interest formula is: FV=PV(1+i)^n Where FV is the final value, PV is the present value, i is the interest rate and n is the time (in years). Are you able to try and work towards an answer now?

OpenStudy (anonymous):

i got 14.1?

OpenStudy (anonymous):

Could you write out the formula after you substituted in the values please? I got a different answer.

OpenStudy (anonymous):

wat did u get and how

OpenStudy (anonymous):

Please write out what you get after you put in the numbers and I'll see where you went wrong.

OpenStudy (anonymous):

jus explain how u set it up and how u got the awnser....

OpenStudy (anonymous):

Use the formula I just gave you. 5000(1+0.0625)^n=10000 Do you agree so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So divide through by 5000 to get (1.0625)^n=2 Then use logs to find n. Are you able to finish the question?

OpenStudy (anonymous):

logs?

OpenStudy (anonymous):

Logarithms: \[\log_{1.0625} 2=n\]

OpenStudy (anonymous):

11.4?

OpenStudy (anonymous):

Exactly. Do you understand the method or is there something I could help explain?

OpenStudy (anonymous):

i got it i jus hat trouble figuring out how to input it on my calculator